# Calibration Attacks: A Comprehensive Study of Adversarial Attacks on Model Confidence

### v2

@@ Abstract
In this work, we highlight and perform a comprehensive study on calibration attacks, a form of adversarial attacks that aim to trap victim models to be heavily miscalibrated without altering their predicted labels, hence endangering the trustworthiness of the models and follow-up decision making based on their confidence. We propose four typical forms of calibration attacks: underconfidence, overconfidence, maximum miscalibration, and random confidence attacks, conducted in both the black-box and white-box setups. We demonstrate that the attacks are highly effective on both convolutional and attention-based models: with a small number of queries, they seriously skew confidence without changing the predictive performance. Given the potential danger, we further investigate the effectiveness of a wide range of adversarial defence and recalibration methods, including our proposed defences specifically designed for calibration attacks to mitigate the harm. From the ECE and KS scores, we observe that there are still significant limitations in handling calibration attacks. To the best of our knowledge, this is the first dedicated study that provides a comprehensive investigation on calibration-focused attacks. We hope this study helps attract more attention to these types of attacks and hence hamper their potential serious damages. To this end, this work also provides detailed analyses to understand the characteristics of the attacks.
@@ 1 Introduction
While recent machine learning models have significantly improved the state-of-the-art performance on a wide range of tasks (Bengio et al., 2021; Vaswani et al., 2017; LeCun et al., 2015; Krizhevsky et al., 2012), these models are often vulnerable and easily deceived by perturbed input (Ren et al., 2020). Adversarial attacks (Ren et al., 2020) have been shown to be a crucial tool to reveal the susceptibility of victim models  (Ibitoye et al., 2019; Zimmermann et al., 2022; Xiao et al., 2023). In the classic setup, adversarial examples are generated by introducing an imperceptible modification to an original datapoint to cause misclassification, where the focus is on trapping victim models to make incorrect predictions.
In this paper, we highlight and provide a comprehensive study on a different type of threats, which we call calibration attacks. The attacks focus on the victim models’ confidence scores without modifying their predicted labels, hence endangering any follow-up decision-making that is based on the victim models’ confidence. We propose to conduct four forms of calibration attacks: underconfidence, overconfidence, maximum miscalibration, and random confidence attacks, which can seriously skew confidence and cause heavy miscalibration, as demonstrated in the reliability diagrams in Figure 1. As we will show and discuss in our study, calibration attacks are insidious and hard to detect. The intrinsic harm is that on the surface the models appear to still make correct decisions, but the level of miscalibration could make the models’ decisions malicious for downstream tasks.
Our specific studies consist of four forms of attacks, span over black-box and white-box setups, attack typical convolutional and attention-based models, and investigate both attack and defence methods. In summary, our main contributions are as follows.
@@
  • We propose to perform four typical forms of calibration attacks and demonstrate their effectiveness and danger to victim models from different perspectives. Detailed insights are provided to understand the characteristics of the attacks and the vulnerability of victim models.
@@ 2 Related Work
Among the two typical types of calibration methods, post calibration is applied directly to the predictions of fully trained models at test time, which include classical approaches such as temperature scaling (Guo et al., 2017), Platt scaling (Platt, 1999), isotonic regression (Zadrozny & Elkan, 2002), and histogram binning (Zadrozny & Elkan, 2001). Training-based approaches, however, often add bias to help calibrate a model during training (Zhang et al., 2018; Thulasidasan et al., 2019; Kumar et al., 2018; Tomani & Buettner, 2021). In our work we investigate a diverse range of calibration methods against calibration attacks to real the limitations that require them to be overhauled to deal with attacks, including the vulnerability of models on the convolutional architectures (Guo et al., 2017; Minderer et al., 2021) and Transformer-based frameworks (Dosovitskiy et al., 2021). (Refer to Appendix A for a more detailed summary of related work.)
@@
Adversarial attacks include black-box and white-box approaches. The former  (Carlini & Wagner, 2017) assume less information about victim models. Many methods are based on gradient estimation through querying the models and finding the finite differences (Bhagoji et al., 2018). White-box attacks often have access to the full details of a victim model such as model architectures and gradients (Goodfellow et al., 2015). Adversarial training and defence approaches have been introduced to improve the robustness of victim models (Stutz et al., 2020; Chen et al., 2022; Qin et al., 2021; Patel et al., 2021; Dhillon et al., 2018). Related to these are works on generating certified robustness guarantees Kumar et al. (2020), where certified radii are generated for the predicted confidence of a smoothed classifier. The research in Emde et al. (2023) further extends this by certifying calibration through generating the worst-case bounds of calibration error, and discuss the importance studying attacks targeting calibration. In this work we propose two defence models against calibration attacks.
@@ 3.1 Objective of Calibration Attacks
Calibration attacks aim to generate adversarial examples to optimize a predefined miscalibration function $\mathcal{M}(\tilde{\mathbf{x}}_{i},k)$ for an adversarial example $\tilde{\mathbf{x}}_{i}$ and the predicted class $k$. As will be detailed below, we propose four forms of calibration attacks where $\mathcal{M}(\tilde{\mathbf{x}}_{i},k)$ takes different implementations. Following the conventional notations, an adversarial example $\tilde{\mathbf{x}}_{i}$ is created by adding noise $\bm{\delta}$ to an input $\mathbf{x}_{i}$: $\tilde{\mathbf{x}}_{i}=\mathbf{x}_{i}+\bm{\delta}$, bounded by $\epsilon$ in a $l_{m}$-ball:
@@ 3.2 Four Forms of Calibration Attacks
$\mathcal{M}_{\textit{UCA}}{(\tilde{\mathbf{x}}_{i},k)}$ $=\hat{p}_{k}(\tilde{\mathbf{x}}_{i})-\max_{j\neq k}\hat{p}_{j}(\tilde{\mathbf{x}}_{i}),$ (3)
@@
As discussed earlier in introduction, calibration attacks focus on attacking the confidence of victim models but not altering the predicted labels, which is constrained by Eq. 5. By minimizing the loss $\mathcal{M}_{\textit{UCA}}$ or $\mathcal{M}_{\textit{OCA}}$, the attack models maximize calibration errors in these two setups respectively, with the corresponding adversarial examples $\bm{\mathbf{\tilde{X}}}$ :
@@
Algorithm 1 depicts an overview of calibration attacks, including UCA and OCA, but also MMA and RCA that we will introduce below. Unlike the conventional attacks that modify the predicted labels and focus mainly on correctly classified examples, calibration attacks modify confidence and focus on both the originally correctly classified and misclassified instances. As detailed later in Section 4, we will implement calibration attacks in popular black-box  (Andriushchenko et al., 2020) and white-box frameworks (Madry et al., 2018).
@@
20:     $\bm{\delta}=FindPerturb\;(\tilde{\mathbf{x}}_{i},\epsilon)$ // Find perturbation based on an attack algorithm (e.g., Square Attack) and bound.
@@
Maximum Miscalibration Attacks (MMA). We propose MMA with the aim of exploring and understanding more serious scenarios of miscalibration. The main principle of MMA is the aim to perturb (i) all incorrectly classified datapoints to have zero accuracy and 100% confidence, and (ii) all correctly classified attacks to have 100% accuracy but the minimum possible confidence. MMA is a combination of overconfidence and underconfidence attack, as shown in Algorithm 1. Proposition 3.1 below states the property of MMA in terms of the oracle (upper-bound) ECE score that can be achieved in theory, with the proof provided in Appendix B.
@@
Assume $q$ is the accuracy of a K-way classifier $\mathcal{F}$ on the dataset $\mathcal{D}=\{\langle\mathbf{x}_{n},y_{n}\rangle\}^{N}_{n=1}$. The Maximum Misinformation Attack (MMA) maximizes the expected calibration error (ECE). The upper bound of ECE that can be achieved by MMA is $1-q/K$.
Random Confidence Attacks (RCA). We propose to perform RCA to decouple the victim model’s confidence scores and predictive labels, in which the confidence scores produced by the attack model are randomized. RCA is performed by choosing a random target confidence score for each input, and then, depending on the original confidence score, running the corresponding underconfidence or overconfidence attacks to produce the target confidence score. Although RCA is theoretically less effective than MMA, it is less predictable, because unlike the other three types of attacks, RCA does not produce adetermined direction (i.e., under or overconfidence) of attacks for a given input. The produced confidence scores are often less extreme and more reasonable-looking, but largely meaningless due to randomization. Note that in addition to MMA and RCA, one may design other approaches to combine UCA with OCA, but MMA and RCA represent the most typical composition.
@@ 4.1 Experimental Setup
Implementation Deails. As discussed earlier in this paper, calibration attacks can be built on different adversarial attack frameworks. In our implementation we use Square Attack (SA)  (Andriushchenko et al., 2020), which is one of the most popular black-box approaches and is highly effective. SA achieves state-of-the-art (SOTA) performance in terms of query efficiency and success rate, even outperforming some white-box methods. Our white-box calibration attacks are based on the popular Projected Gradient Descent (PGD) framework (Madry et al., 2018). The implementation details such as hyperparameters are discussed in Appendix C.
Models. We include both convolutional (ResNet (He et al., 2016)) and non-convolutional attention-based models (Vision Transformer (ViT) (Dosovitskiy et al., 2021)) in our study. Details can be found in Appendices C and D.
@@
Metrics. Two calibration metrics are used in our experiments: the stardard Expected Calibration Error (ECE) (Pakdaman Naeini et al., 2015) and the recent Kolmogorov-Smirnov Calibration Error (KS error) (Gupta et al., 2021). In addition, we evaluate attacks’ efficiency using the average and median number of queries for the attack to complete (Andriushchenko et al., 2020). Average confidence of predictions is also leveraged to judge the degree that the confidence scores are affected. (See Appendix C for details.)
Detailed Attack Settings. Descriptions of other attack settings (e.g., $l_{\infty}$, $l_{2}$ and iterations) are in Appendix Cand D.
4.2 Overall Performance
Table 1 depicts the overall performance of black-box attacks under the $l_{\infty}$ norm. Additional results of $l_{2}$ attacks are included in Table 6 in Appendix G.9. The experiments show that the attacks are highly effective. UCA, MMA and RCA can bring ECE and KS to very high values. OCA can successfully raise average confidences (the last column in Table 1) to extremely high levels (in many cases, close to 100%), but given the original high accuracy of the victim models, increasing confidence levels will not have a drastic effect on calibration error. (OCA will be more harmful on less accurate models.) For the MMA, the theoretical highest levels of miscalibration are not reached due to the limited number of iterations that we run the attacks. Regarding different architectures, the attention-based ViT models are seen to be more miscalibrated compared to the convolution-based ResNet models. The white-box results are included in Appendix G.1, which show a similar trend. Again, Figure 1 in the introduction section demonstrates the attack effects using calibration diagrams, visually showing the severity of miscalibration.
The general trend of the $l_{2}$ calibration attacks (Appendix G.9) is similar to that of the $l_{\infty}$ attacks, but the latter are found to be more effective in generating more significant miscalibration. Hence in the remainder of the paper, we focus on the $l_{\infty}$ attacks unless otherwise specified.
@@ 4.3 Detection Difficulty Analysis
This section shows calibration attacks are also difficult to detect. To investigate this, we run the popular adversarial attack detection methods on the attacks against ResNet-50: Local Intrinsic Dimensionality (LID) (Ma et al., 2018), Mahalanobis Distance (MD) (Lee et al., 2018), and SpectralDefense (Harder et al., 2021). The details behind the settings for each detection method can be found in Appendix F. Table 2 depicts the main results of the effectiveness of the detectors in terms of Area Under the Curve (AUC) and Detection Accuracy, under different types of calibration attacks and using both the white-box and black-box approaches. We can see that there are consistent decreases in detection performances, particularly in SA, where the decreases are often more than 20%. The existing detection methods are shown to be less reliable for calibration attacks.

### v3

@@ Abstract
In this work, we highlight and perform a comprehensive study on calibration attacks, a form of adversarial attacks that aim to trap victim models to be heavily miscalibrated without altering their predicted labels, hence endangering the trustworthiness of the models and follow-up decision making based on their confidence. We propose four typical forms of calibration attacks: underconfidence, overconfidence, maximum miscalibration, and random confidence attacks, conducted in both black-box and white-box setups. We demonstrate that the attacks are highly effective on both convolutional and attention-based models: with a small number of queries, they seriously skew confidence without changing the predictive performance. Given the potential danger, we further investigate the effectiveness of a wide range of adversarial defence and recalibration methods, including our proposed defences specifically designed for calibration attacks to mitigate the harm. From the ECE and KS scores, we observe that there are still significant limitations in handling calibration attacks. To the best of our knowledge, this is the first dedicated study that provides a comprehensive investigation on calibration-focused attacks. We hope this study helps attract more attention to these types of attacks and hence hamper their potential serious damages. To this end, this work also provides detailed analyses to understand the characteristics of the attacks.  [^1] Our code is available at https://github.com/PhenetOs/CalibrationAttack
@@ 1 Introduction
While recent machine learning models have significantly improved the state-of-the-art performance on a wide range of tasks (Bengio et al., 2021; Vaswani et al., 2017; LeCun et al., 2015; Krizhevsky et al., 2012), these models are often vulnerable and easily deceived by perturbed inputs (Ren et al., 2020). Adversarial attacks (Ren et al., 2020) have been shown to be a crucial tool to reveal the susceptibility of victim models  (Ibitoye et al., 2019; Zimmermann et al., 2022; Xiao et al., 2023). In the classic setup, adversarial examples are generated by introducing imperceptible modifications to an original datapoint to cause misclassification, where the focus is on trapping victim models to make incorrect predictions.
In this paper, we highlight and provide a comprehensive study on a different type of threat, which we call calibration attacks. These attacks focus on manipulating the victim models’ confidence scores without modifying their predicted labels, hence endangering any follow-up decision-making that is based on the victim models’ confidence. We propose to conduct four forms of calibration attacks: underconfidence, overconfidence, maximum miscalibration, and random confidence attacks, which can seriously skew confidence and cause heavy miscalibration, as demonstrated in the reliability diagrams in Figure 1. As we will show and discuss in our study, calibration attacks are insidious and hard to detect. The intrinsic harm is that on the surface the models appear to still make correct decisions, but the level of miscalibration could make the models’ decisions malicious for downstream tasks.
Our specific studies consist of four forms of attacks, spanning black-box and white-box setups. We attack typical convolutional and attention-based models, and investigate both attack and defence methods.
In summary, our main contributions are as follows.
@@
  • We propose four typical forms of calibration attacks and demonstrate their effectiveness and danger to victim models from different perspectives. Detailed insights are provided to understand the characteristics of the attacks and the vulnerability of victim models.
@@ 2 Related Work
Among the two typical types of calibration methods, post-calibration methods are applied directly to the predictions of fully trained models at test time, which include classical approaches such as temperature scaling (Guo et al., 2017), Platt scaling (Platt, 1999), isotonic regression (Zadrozny & Elkan, 2002), and histogram binning (Zadrozny & Elkan, 2001). Training-based approaches, however, often add bias to help calibrate a model during training (Zhang et al., 2018; Thulasidasan et al., 2019; Kumar et al., 2018; Tomani & Buettner, 2021). We investigate a diverse range of calibration methods against calibration attacks to reveal the limitations that require them to be overhauled to deal with attacks, examining the vulnerability of models similar to those on the convolutional architectures (Guo et al., 2017; Minderer et al., 2021) and Transformer-based frameworks (Dosovitskiy et al., 2021). (Refer to Appendix A for a more detailed summary of related work.)
@@
Adversarial attacks are divided into black-box and white-box approaches. The former  (Carlini & Wagner, 2017) assume less information about victim models. Many methods are based on gradient estimation through querying the models and finding the finite differences (Bhagoji et al., 2018). White-box attacks often have access to the full details of a victim model such as its architecture and gradients (Goodfellow et al., 2015). Adversarial training and defence approaches have been introduced to improve the robustness of victim models (Chen et al., 2022; Qin et al., 2021; Patel et al., 2021; Dhillon et al., 2018). Related to these are works on generating certified robustness guarantees Kumar et al. (2020), where certified radii are generated for the predicted confidence of a smoothed classifier. The research in Emde et al. (2023) further extends this by certifying calibration through generating the worst-case bounds of calibration error, and discuss the importance studying attacks targeting calibration. Stutz et al. (2020) develop a modified adversarial training objective which focuses on biasing adversarial samples to be low-confidence predictions that can be easily filtered. Although this causes adversarial examples to be low confidence, it does not specifically target making low or high confidence adversarial examples through perturbations as a form of attack. In this work, we propose two defence models against calibration attacks.
@@ 3.1 Objective of Calibration Attacks
Calibration attacks aim to generate adversarial examples to optimize a predefined miscalibration function $\mathcal{M}(\tilde{\mathbf{x}}_{i},k)$ for an adversarial example $\tilde{\mathbf{x}}_{i}$ and the predicted class $k$. As will be detailed below, we propose four forms of calibration attacks where $\mathcal{M}(\tilde{\mathbf{x}}_{i},k)$ takes different implementations. Following conventional notation, an adversarial example $\tilde{\mathbf{x}}_{i}$ is created by adding noise $\bm{\delta}$ to an input $\mathbf{x}_{i}$: $\tilde{\mathbf{x}}_{i}=\mathbf{x}_{i}+\bm{\delta}$, bounded by $\epsilon$ in a $l_{m}$-ball:
@@ 3.2 Four Forms of Calibration Attacks
$\mathcal{M}_{\textit{UCA}}{(\tilde{\mathbf{x}}_{i},k)}$ $=\hat{p}_{k}(\tilde{\mathbf{x}}_{i})-\max_{\rm{j\neq k}}\hat{p}_{j}(\tilde{\mathbf{x}}_{i}),$ (3)
@@
Calibration attacks focus on attacking the confidence of victim models without altering the predicted labels, which is satisfied by the constraint in Eq. 5.
By minimizing the loss $\mathcal{M}_{\textit{UCA}}$ or $\mathcal{M}_{\textit{OCA}}$, the attack models maximize calibration errors in these two setups respectively, with the corresponding adversarial examples $\bm{\mathbf{\tilde{X}}}$ :
@@
Algorithm 1 depicts an overview of calibration attacks, including UCA and OCA, but also MMA and RCA that we will introduce below. Unlike conventional attacks that modify the predicted labels and focus mainly on correctly classified examples, calibration attacks modify confidence and focus on both the originally correctly classified and misclassified instances. As detailed later in Section 4, we will implement calibration attacks in popular black-box  (Andriushchenko et al., 2020) and white-box frameworks (Madry et al., 2018).
@@
20:     $\bm{\delta}=FindPerturb\;(\tilde{\mathbf{x}}_{i},\epsilon)$ // Find perturbation based on attack algorithm (e.g., Square Attack) and bound.
@@
Maximum Miscalibration Attacks (MMA). We propose MMA with the aim of exploring and understanding more serious scenarios of miscalibration. The main principle of MMA is to perturb a set of inputs such that (i) all incorrectly classified datapoints (a set where the model has zero accuracy) have 100% confidence, and (ii) all correctly classified datapoints (a set where the model is 100% accurate) have the minimum possible confidence. MMA uses a combination of OCA and UCA to accomplish this, as shown in Algorithm 1. Proposition 3.1 below states the property of MMA in terms of the oracle (upper-bound) ECE score that can be achieved in theory, with the proof provided in Appendix C.
@@
Assume $q$ is the accuracy of a K-way classifier $\mathcal{F}$ on the dataset $\mathcal{D}=\{\langle\mathbf{x}_{n},y_{n}\rangle\}^{N}_{n=1}$. The Maximum Miscalibration Attack (MMA) maximizes the expected calibration error (ECE). The upper bound of ECE that can be achieved by MMA is $1-q/K$.
Random Confidence Attacks (RCA). We propose to perform RCA to decouple the victim model’s confidence scores and predictive labels, in which the confidence scores produced by the attack model are randomized. RCA is performed by choosing a random target confidence score for each input, and then, depending on the original confidence score, running the corresponding underconfidence or overconfidence attacks to produce the target confidence score. Although RCA is theoretically less effective than MMA, it is less predictable, because unlike the other three types of attacks, RCA does not produce a predetermined direction (i.e., under- or overconfidence) of attacks for a given input. The produced confidence scores are often less extreme and more reasonable-looking, but largely meaningless due to randomization. Note that in addition to MMA and RCA, one may design other approaches to combine UCA with OCA, but MMA and RCA represent the two most typical compositions that create very harmful patterns of miscalibration.
@@ 3.4 Discussion on the Importance of Remaining Well Calibrated Under the Attacks
In addition to the discussion of the technical details of calibration attacks for the victim models and defences, it is important to reemphasize why maintaining good calibration under such attacks is critical. Well-calibrated models are a crucial component for trustworthy complex systems. In terms of specific real-world applications, skewed confidence scores or uncertainties have been shown to have dramatic impacts on downstream tasks and human decision-making, and previous studies such as those of Penso et al. (2024); Jiang et al. (2012); Kompa et al. (2021) have shown the effect of miscalibration on down-stream authority mechanisms in the context of areas like medical imaging. Specific use-cases can also be found in child maltreatment hotline screening models De-Arteaga et al. (2020), legal domain Bernsohn et al. (2024), and anomalous behaviour detectors Shashanka et al. (2016) where heavily miscalibrated outputs present a failure point to methods.
Moreover, prior works (Zhang et al., 2020; Rechkemmer & Yin, 2022) have examined the role of skewed confidence affecting the level of trust humans have in a system, which influences applications where humans and AI decision-making occurs symbiotically. These works demonstrate that when models produce high confidence scores, users are more prone to rely on the AI for decision-making. Dhuliawala et al. (2023) conduct user studies that show that confidently incorrect predictions lead to over reliance and heavily undermine user trust in human-AI collaboration settings, requiring only a few incorrect instances with inaccurate confidence estimates to damage user trust and performance. Please refer to Appendix B for detailed discussion on downstream impacts.
@@ 4.1 Experimental Setup
Implementation Details. As discussed earlier, calibration attacks can be built on different adversarial attack frameworks. In our implementation, we use Square Attack (SA)  (Andriushchenko et al., 2020), which is one of the most popular black-box approaches and is highly effective. SA achieves state-of-the-art (SOTA) performance in terms of query efficiency and success rate, even outperforming some white-box methods. Our white-box calibration attacks are based on the popular Projected Gradient Descent (PGD) framework (Madry et al., 2018). The implementation details such as hyperparameters are discussed in Appendix D.
Models. We include both convolutional (ResNet (He et al., 2016)) and non-convolutional attention-based models (Vision Transformer (ViT) (Dosovitskiy et al., 2021)) in our study. Details can be found in Appendices D and E.
@@
Metrics. Two calibration metrics are used in our experiments: the standard Expected Calibration Error (ECE) (Pakdaman Naeini et al., 2015) and the recent Kolmogorov-Smirnov Calibration Error (KS error) (Gupta et al., 2021). In addition, we evaluate attacks’ efficiency using the average and median number of queries for the attack to complete (Andriushchenko et al., 2020). Average confidence of predictions is also leveraged to judge the degree that the confidence scores are affected. (See Appendix D for details.)
Same as in previous works (Guo et al., 2017; Kumar et al., 2018; Minderer et al., 2021), we focus on intrinsic evaluation metrics in this research, which isolate our analysis from downstream tasks to abstain from task-specific conclusions, particularly considering the downstream tasks here are very diverse. In addition, the harm from miscalibration in downstream applications has been well studied in prior work, as detailed in Appendix B.
Detailed Attack Settings. Descriptions of other attack settings (e.g., $l_{\infty}$, $l_{2}$ and iterations) are in Appendix Dand E.
4.2 Overall Performance
@@
Table 1 depicts the overall performance of black-box attacks under the $l_{\infty}$ norm. Additional results of $l_{2}$ attacks are included in Table 9 in Appendix H.10. The experiments show that the attacks are highly effective. UCA, MMA and RCA can bring ECE and KS to very high values. OCA can successfully raise average confidences (the last column in Table 1) to extremely high levels (in many cases, close to 100%), but given the original high accuracy of the victim models, increasing confidence levels will not have a drastic effect on calibration error. (OCA will be more harmful on less accurate models.) For the MMA, the theoretical highest levels of miscalibration are not reached due to the limited number of iterations that we run the attacks. Regarding different architectures, the attention-based ViT models are seen to be more miscalibrated compared to the convolution-based ResNet models. The white-box results are included in Appendix H.1. While ResNet and ViT are the most representative frameworks for convolution and attention-based models respectively, in Appendix H.8 we include additional experiments on the Swin Transformer (Liu et al., 2021), which is a famous extension of ViT. The results show similar trends as those presented in Table 1. Again, Figure 1 in the introduction section demonstrates the attack effects using calibration diagrams, visually showing the severity of miscalibration.
The general trend of the $l_{2}$ calibration attacks (Appendix H.10) is similar to that of the $l_{\infty}$ attacks, but the latter are found to be more effective in generating more significant miscalibration. Hence, in the remainder of the paper, we focus on the $l_{\infty}$ attacks unless otherwise specified.
@@ 4.3 Detection Difficulty Analysis
This section shows calibration attacks are also difficult to detect. To investigate this, we run the popular adversarial attack detection methods on the attacks against ResNet-50: Local Intrinsic Dimensionality (LID) (Ma et al., 2018), Mahalanobis Distance (MD) (Lee et al., 2018), and SpectralDefense (Harder et al., 2021). The details behind the settings for each detection method can be found in Appendix G. Table 2 depicts the main results of the effectiveness of the detectors in terms of Area Under the Curve (AUC) and Detection Accuracy, under different types of calibration attacks and using both the white-box and black-box approaches. We can see that there are consistent decreases in detection performances, particularly in SA, where the decreases are often more than 20%. The existing detection methods are shown to be less reliable for calibration attacks.